Solve.
w^4 - 4w^2 - 2 = 0
What is the solution? w = _____

Answer by stanbon(75887)

w^4 - 4w^2 - 2 = 0
w^2 = [4 +- sqrt(16-4*1*-2)]/2
w^2 = [4 +- sqrt(24)]/2
w^2 = [4 +- 2sqrt(6)]/2
w^2 = [2 +- sqrt6]
w^2 = 2+sqrt6 or x^2 = 2-sqrt6
w = sqrt(2+sqrt6) or w = -(2+sqrt6) or w=sqrt(2-sqrt6) or w=-sqrt(2-sqrt6)
===================
Cheers,
Stan H.